### v1

# Effects of personality steering on cooperative behavior in Large Language Model agents

Abstract
Large language models (LLMs) are increasingly used as autonomous agents in strategic and social interactions. Although recent studies suggest that assigning personality traits to LLMs can influence their behavior, how personality steering affects cooperation under controlled conditions remains unclear. In this study, we examine the effects of personality steering on cooperative behavior in LLM agents using repeated Prisoner’s Dilemma games. Based on the Big Five framework, we first measure basic personality profiles of three models, GPT-3.5-turbo, GPT-4o, and GPT-5, using the Big Five Inventory. We then compare behavior under baseline and personality-informed conditions, and further analyze the effects of independently manipulating each personality dimension to extreme values. Our results show that agreeableness is the dominant factor promoting cooperation across all models, while other personality traits have limited impact. Explicit personality information increases cooperation but can also raise vulnerability to exploitation, particularly in earlier-generation models. In contrast, later-generation models exhibit more selective cooperation. These findings indicate that personality steering acts as a behavioral bias rather than a deterministic control mechanism.
keywords:
Large Language Model, Personality steering, Cooperative behavior, Repeated Prisoner’s Dilemma
1 Introduction
Multi-agent systems have long been studied and utilized to make decisions on behalf of humans. Traditional multi-agent systems were designed to achieve rational or cooperative behavior under specific environments, based on limited information-processing capabilities and predefined rules. Typical applications include price negotiations and contract adjustments, where multiple agents seek agreement while pursuing conflicting interests [1]; resource allocation scenarios, where competition and cooperation coexist over limited resources [2]; and marketplaces or auctions, where seller and buyer agents strategically propose prices [3, 1]. In recent years, the emergence of Large Language Models (LLMs) has substantially expanded the decision-making capabilities of agents. By employing LLMs as agents, flexible interpretation of information and reasoning through natural language become possible, enabling decision-making even in complex and context-dependent social situations that were previously difficult to handle. As a result, LLM-based multi-agent systems have demonstrated new potential in social interactions involving multiple stakeholders and in settings that require human-like judgment [4, 5, 6, 7].
However, this shift in agent capabilities has also introduced new challenges. Interactions among LLM-based agents can be unpredictable [4] and may involve the risk of escalating unintended conflicts [8]. In strategic environments, optimal decision-making requires not only the maximization of immediate rewards but also careful consideration of relationships with other agents and the formation of long-term cooperative relationships. Such instability highlights the need for new approaches to guide and control the behavior and decision-making of LLMs in strategic settings. As a potential solution to this challenge, the psychological framework of personality has attracted increasing attention. In personality psychology, various theories have been proposed to explain individual behavioral patterns, among which the Big Five Personality Traits [9] are widely accepted as a theoretical framework that describes personality along five major dimensions: (1) Extraversion, which reflects sociability, activity level, and the tendency to seek stimulation; (2) Agreeableness, which captures empathy, cooperativeness, and tolerance toward others; (3) Conscientiousness, which represents responsibility, organization, and motivation toward goal achievement; (4) Neuroticism, which reflects emotional instability and sensitivity to stress; and (5) Openness to Experience, which captures openness to new experiences and ideas, as well as creativity. In human psychological research, individuals with high agreeableness tend to exhibit stronger cooperative behavior, while also being more vulnerable to exploitation [10]. Moreover, in strategic situations such as repeated Prisoner’s Dilemma games, individuals with higher agreeableness are more likely to cooperate in the early stages of interaction [11].
At the same time, because the Big Five framework provides well-established quantitative measurement methods, it is also well suited as a basis for analyzing personality traits in LLMs, suggesting that such traits may influence decision-making in these models. Indeed, a previous study has reported that increasing traits such as agreeableness and conscientiousness in LLM agents promotes cooperative behavior, while simultaneously increasing vulnerability to exploitation by other agents, leading to lower overall payoffs-patterns that closely resemble those observed in humans [12]. However, this prior study has several important limitations. First, it typically begins by assigning or manipulating personality traits in LLMs without first quantitatively measuring the basic personality profiles inherently exhibited by the models, making it difficult to assess how personality interventions alter behavior relative to a model’s original tendencies. Second, because personality traits are often not specified in quantitative terms, the strength of personality manipulations cannot be systematically compared, which limits reproducibility and hinders meaningful comparisons across studies.
To address these limitations, we investigate the relationship between personality traits and cooperative behavior in LLM agents under quantitatively controlled conditions using the Big Five Personality Traits framework. Specifically, we examine three objectives. First, we quantitatively measure the basic personality profiles inherently exhibited by different LLMs based on the Big Five framework, thereby clarifying model-specific characteristics and differences across model generations. Second, we examine how LLM agents behave in strategic settings such as repeated Prisoner’s Dilemma games when their measured personality traits are explicitly provided through prompts. Third, we analyze how cooperative behavior changes when each personality dimension is independently manipulated to extreme values while holding the remaining dimensions constant. Through these analyses, we aim to clarify how personality steering influences cooperative behavior in LLM agents and to provide a controlled framework for understanding the role of personality traits in LLM decision-making.
2 Methods
In this study, we investigate the relationship between personality traits and cooperative behavior in LLM agents using three widely used models developed by OpenAI: GPT-3.5-turbo, GPT-4o, and GPT-5. We conducted the experiments in three stages. First, we quantitatively measured the basic personality profiles of each model based on the Big Five Personality Traits using the Big Five Inventory (BFI-44) (Experiment 1). Second, we examined how LLM agents behave in strategic settings by having them play repeated Prisoner’s Dilemma games under two conditions: a baseline condition without explicit personality information and a condition in which the measured personality traits obtained in Experiment 1 were explicitly provided via prompts (Experiment 2). Third, we analyzed the effects of personality manipulation by independently setting each Big Five trait to extreme values while keeping the remaining traits constant, in order to assess which personality dimensions most strongly influenced cooperative behavior (Experiment 3).
2.1 Experiment 1: Measuring basic personality traits using the Big Five
In Experiment 1, we first measure the personality traits of each model based on the Big Five Personality Traits framework [9]. To quantitatively assess these personality traits in LLMs, we employed the Big Five Inventory (BFI), one of the most widely used psychometric instruments for measuring personality traits [9]. The BFI is a self-report questionnaire developed based on the Big Five theory and has been validated across diverse cultures and languages, demonstrating high reliability and validity. In this study, we used the BFI-44, a shortened version consisting of 44 items, with 8–10 items assigned to each personality dimension. Each item is rated on a five-point Likert scale ranging from 1 (“strongly disagree”) to 5 (“strongly agree”).
For each model, the 44 questionnaire items were presented sequentially, with instructions prompting the model to respond to each item using a numerical value from 1 to 5 (see Appendix A for the full prompt). We adopt the BFI prompt developed by Jiang et al. [13]. If a response did not conform to the required format (e.g., contains non-numeric characters), the question was presented again until a valid numerical response was obtained. After collecting responses to all items, scores for each personality dimension were computed following the standard BFI scoring procedure. Because LLM outputs involve stochasticity, we repeated the personality measurement five times for each model and used the average scores across the five runs to improve score stability. Through this procedure, we obtained quantitative and reproducible basic personality profiles for each LLM.
2.2 Experiment 2: Behavioral differences with explicit personality information
In Experiment 2, we examine how LLM agents behave in strategic environments by having them play repeated Prisoner’s Dilemma (RPD) games. We considered two experimental conditions. In the baseline condition, no explicit personality information was provided to the LLMs. In the personality-informed condition, the personality traits measured in Experiment 1 were explicitly provided to the LLMs via prompts. By comparing these two conditions, we assessed whether and how explicitly supplied personality information influenced cooperative behavior.
The Prisoner’s Dilemma is a canonical game-theoretic model that captures social dilemmas arising from the tension between cooperation and defection. In each round, two players simultaneously choose either to cooperate (C) or defect (D). Payoffs are determined by the joint action of both players according to a payoff matrix satisfying $T>R>P>S$ , where $T$ denotes the temptation to defect, $R$ the reward for mutual cooperation, $P$ the punishment for mutual defection, and $S$ the payoff received by a cooperator when the opponent defects. In this study, we used the standard payoff values $T=5$, $R=3$, $P=1$, and $S=0$.
In the RPD game, the same pair of players interacts repeatedly, and it is assumed that each player can observe and remember past actions and outcomes. This repeated structure allows players to condition their current decisions on previous interactions, thereby enabling the emergence of cooperative behavior. Both theoretical [14, 15, 16, 17, 18, 19, 20] and experimental [21, 22, 23, 24, 25] studies have demonstrated that cooperation can be sustained in repeated Prisoner’s Dilemma games under appropriate conditions.
We first conducted the RPD experiments under the baseline condition to characterize the intrinsic cooperative tendencies of each LLM. In this condition, the prompt included only the rules and objectives of the RPD game and did not contain any information related to personality traits. The prompt was constructed based on that proposed by Fontana et al. [26]. Each LLM agent played 10 independent trials, each consisting of 10 rounds of the RPD game. Trials were independent, and outcomes from previous trials did not affect subsequent trials.
In each trial, the LLM agent played against a set of fixed opponent strategies that were hard-coded and commonly used in studies of the RPD game. These strategies include: Always Cooperate (ALLC), which selects cooperation in every round; Always Defect (ALLD), which selects defection in every round; Random (RANDOM), which randomly chooses between cooperation and defection in each round; Tit-for-Tat (TFT), which cooperates in the first round and then replicates the opponent’s previous action; and Grim Trigger (GRIM), which cooperates initially but switches permanently to defection after the opponent defects once. Together, these strategies cover a broad range of cooperative and non-cooperative behaviors.
In the personality-informed condition, we repeated the same RPD experiments while explicitly providing each LLM with its measured Big Five personality scores obtained in Experiment 1. The prompt specifies the personality traits in the form: “Your personality traits are as follows: Openness (O), Conscientiousness (C), Extraversion (E), Agreeableness (A), and Neuroticism (N),” followed by the corresponding numerical scores (see Appendix for prompt details). Under this condition, the LLM was instructed to make decisions while being aware of its own personality traits.
For each experimental condition, we computed two evaluation metrics. The average cooperation rate is defined as the proportion of cooperative choices across all rounds, indicating the degree of cooperativeness exhibited by the LLM. The average cumulative payoff is defined as the mean total payoff accumulated across all rounds, reflecting the overall performance of the LLM. By comparing these metrics between the baseline and personality-informed conditions, we evaluated the behavioral impact of explicitly providing personality information to LLM agents.
2.3 Experiment 3: Effects of extreme personality manipulation
In Experiment 3, we investigate how artificial manipulation of personality traits affects the behavior of LLM agents by modifying the Big Five personality scores obtained in Experiment 1. Specifically, we independently set one personality dimension, Agreeableness, Extraversion, Conscientiousness, Neuroticism, or Openness, to an extreme value of either 1 or 5, while keeping the remaining four dimensions fixed at their measured average values. This procedure allows us to isolate the individual contribution of each personality dimension to cooperative behavior.
In each experimental condition, only one personality trait was manipulated. For example, when manipulating agreeableness, the agreeableness score is set to either 1 (minimum) or 5 (maximum), while the scores for openness, conscientiousness, extraversion, and neuroticism remain unchanged from the values measured in Experiment 1. By applying this manipulation independently to each of the five personality dimensions and considering both extreme values, we obtain a total of ten experimental conditions (five dimensions $\times$ two values).
For each condition, the modified personality scores are explicitly provided to the LLM agents via prompts in the same format as in the personality-informed condition of Experiment 2. The agents then play the RPD games under identical settings, including the same opponent strategies, number of rounds, and number of trials. Using the same evaluation metrics as in Experiment 2, the average cooperation rate and the average cumulative payoff, we assessed how extreme manipulation of individual personality traits influenced cooperative behavior.
2.4 LLM Settings
In this study, we used three large language models developed by OpenAI: GPT-3.5-turbo, GPT-4o, and GPT-5. These models represent successive generations of LLMs, with larger model numbers corresponding to more recent releases.
For GPT-3.5-turbo and GPT-4o, we set the temperature parameter to 0.7. The temperature controls the degree of randomness in model outputs, with lower values producing more deterministic responses and higher values increasing stochasticity. The value of 0.7 is chosen to balance response diversity and consistency, and is consistent with the setting used in prior work [26].
For GPT-5, we conducted experiments with the reasoning effort set to minimal and the verbosity set to low. The reasoning effort parameter controls the amount of computational effort allocated to the model’s internal reasoning process; setting it to minimal simplifies the reasoning process and improves response speed. The verbosity parameter controls the level of detail in the model’s output; setting it to low ensures that only the minimum necessary information is produced.
3 Results
3.1 BFI scores
We first report the results of Experiment 1, in which the personality traits of each LLM were measured using the Big Five Inventory. Table 1 presents the mean scores and standard deviations obtained from five repeated measurements for each model. For comparison, the table also includes human data reported by Srivastava et al. [27], where the original POMP scores were rescaled to the same 1–5 scale used in this study.
Personality trait GPT-3.5-turbo GPT-4o GPT-5 Human
Openness (O) 4.60 (0.14) 4.78 (0.04) 4.78 (0.13) 3.98 (0.66)
Conscientiousness (C) 4.09 (0.05) 4.09 (0.09) 4.78 (0.21) 3.55 (0.73)
Extraversion (E) 3.80 (0.19) 3.00 (0.29) 3.10 (0.27) 3.18 (0.90)
Agreeableness (A) 4.15 (0.19) 4.31 (0.31) 4.24 (0.21) 3.66 (0.72)
Neuroticism (N) 1.88 (0.13) 1.89 (0.29) 2.03 (0.10) 3.04 (0.88)
Table 1: Mean BFI scores and standard deviations (in parentheses) for each model.
Across all models, the five measurements yielded highly consistent scores, with relatively small standard deviations. Specifically, the standard deviations ranged from 0.05 to 0.19 for GPT-3.5-turbo, from 0.04 to 0.31 for GPT-4o, and from 0.10 to 0.27 for GPT-5. These values are substantially smaller than those observed in humans (0.66–0.90), indicating that each LLM exhibits a stable and internally consistent personality profile. This consistency suggests that the measured personality traits reliably characterize each model.
In contrast, the human scores cluster around the mid-range of the scale, with larger standard deviations, reflecting substantial inter-individual variability in human personality traits. Compared to human averages, all LLMs exhibit higher scores in openness, conscientiousness, and agreeableness, as well as lower scores in neuroticism. This pattern suggests that LLMs tend to exhibit personality profiles that are systematically shifted toward higher scores in certain traits compared to humans, potentially reflecting biases toward socially desirable response patterns present in their training data.
Next, we examined the patterns observed across individual personality dimensions. Overall, all models exhibited broadly similar tendencies, characterized by relatively high scores in openness, conscientiousness, and agreeableness, and low scores in neuroticism. This consistency across models suggests that contemporary LLMs share common personality profiles as measured by the BFI framework. A particularly notable result is observed for conscientiousness. While GPT-3.5-turbo and GPT-4o show identical mean scores for conscientiousness (4.09), GPT-5 exhibits a substantially higher score (4.78). One possible explanation for this difference is the technical improvements reported for GPT-5, including a significant reduction in hallucinated responses and enhanced reasoning consistency compared to earlier models [28]. Such improvements may contribute to more reliable, goal-oriented, and internally consistent responses, which are reflected in higher conscientiousness scores.
3.2 Comparison between baseline and personality-informed conditions
We next compare the results of the repeated Prisoner’s Dilemma experiments under the baseline condition and the personality-informed condition. Figure 1 shows the average cooperation rates and average cumulative payoffs for each model, where dashed lines indicate the baseline condition and solid lines indicate the personality-informed condition. The average cumulative payoff is normalized to range between 0 and 1 for ease of comparison across models.
Figure 1: Average cooperation rates (left) and average cumulative payoffs (right) under the baseline condition (dashed) and the personality-informed condition (solid).
Across all models, the personality-informed condition yielded cooperation rates that were comparable to or higher than those observed in the baseline condition. This pattern is consistent with the high agreeableness scores measured in Experiment 1 and suggests that explicitly providing personality information, particularly high agreeableness, promotes cooperative behavior in LLM agents. These results indicate that personality steering can systematically influence agent behavior in strategic interactions.
For GPT-3.5-turbo, cooperation rates against non-cooperative strategies such as ALLD and RANDOM were relatively high even in the baseline condition, compared to GPT-4o and GPT-5. Under the personality-informed condition, cooperation rates against these strategies increased substantially, while average cumulative payoffs decreased. This result indicates that although explicit personality information enhances cooperation, it also increases vulnerability to exploitation by defecting or unpredictable opponents. In contrast, cooperation rates and payoffs against cooperative or reciprocal strategies, ALLC, GRIM TRIGGER, and TFT, remained high and largely unchanged across both conditions.
GPT-4o exhibited lower cooperation rates against ALLD and RANDOM strategies in the baseline condition, indicating more cautious behavior toward non-cooperative opponents. However, under the personality-informed condition, cooperation rates against these strategies increased, accompanied by a reduction in cumulative payoffs. Notably, the increase in cooperation was more pronounced for RANDOM than for ALLD, suggesting that GPT-4o distinguishes between purely defecting strategies and those that allow for potential reciprocal cooperation.
For GPT-5, baseline behavior against ALLD and RANDOM was broadly similar to that of GPT-4o, whereas cooperation rates against ALLC, GRIM TRIGGER, and TFT were slightly lower than those observed for the other models. An inspection of action histories reveals that GPT-5 frequently defected in the final round while maintaining cooperation in earlier rounds, a pattern consistent with end-game optimization in finitely repeated games. Explanations generated by the model often refer to balancing long-term trust-building with payoff maximization in the final round, suggesting a more strategic decision-making process. Notably, such payoff-maximizing and seemingly self-interested behavior is consistent with prior findings showing that models with explicit reasoning capabilities tend to exhibit more self-interested decision-making compared to models without explicit reasoning processes [29]. Under the personality-informed condition, cooperation rates against ALLC, GRIM TRIGGER, and TFT approached 100%, while cumulative payoffs decreased. Cooperation rates against ALLD remained largely unchanged, and cooperation against RANDOM increased only modestly, indicating that GPT-5 limited cooperation with exploitative opponents even when personality information promoted cooperativeness.
Comparing the impact of personality steering across models reveals clear generational differences. For GPT-3.5-turbo and GPT-4o, explicit personality information leads to higher cooperation rates but lower cumulative payoffs, reflecting a classic trade-off between cooperation and susceptibility to exploitation. In contrast, GPT-5 exhibits a more selective response: while cooperation with reciprocal strategies increases, cooperation with non-cooperative strategies remains constrained. This pattern suggests that newer models can incorporate personality steering while maintaining strategic robustness.
3.3 Effects of extreme personality manipulation
Finally, we report the results of Experiment 3, in which we artificially manipulated individual Big Five personality traits to extreme values. Using the average BFI scores obtained under the personality-informed condition in Experiment 1 as a reference, we independently replaced one personality dimension, openness, conscientiousness, extraversion, agreeableness, or neuroticism, with an extreme value of either 1 or 5, while keeping the remaining dimensions fixed. The RPD experiments were then conducted under the same settings as in Experiment 2. Figures 2, 3, and 4 present the average cooperation rates, average cumulative payoffs, and their differences for GPT-3.5-turbo, GPT-4o, and GPT-5, respectively.
For GPT-3.5-turbo, manipulation of agreeableness had the strongest impact on cooperative behavior, followed by extraversion. When agreeableness was set to 1 (A-1), cooperation rates were consistently lower than those observed under other personality manipulations across all opponent strategies, resulting in a large negative difference in average cooperation. In contrast, when agreeableness was set to 5 (A-5), cooperation rates were comparable to those observed under other personality conditions, yielding relatively small differences.
Figure 2: Average cooperation rates and average cumulative payoffs for GPT-3.5-turbo under different personality score settings. The top row corresponds to setting the manipulated dimension to 1, the middle row to 5, and the bottom row shows the difference (5-1).
A similar but weaker pattern is observed for extraversion. Under the E-1 condition, cooperation rates against ALLD and RANDOM strategies were slightly lower than those under other personality manipulations, which contributes to observable differences in average cooperation. When extraversion is set to 5 (E-5), however, cooperation rates show little deviation from other conditions.
For GPT-4o, agreeableness again emerged as the dominant factor influencing cooperative behavior, whereas manipulation of other personality traits had minimal effects. Under the A-1 condition, cooperation rates droped to 0% across all opponent strategies, indicating consistently defecting behavior. Under the A-5 condition, average cooperation rates increased to levels comparable to or higher than those observed under other personality manipulations. However, cooperation rates against ALLD and RANDOM strategies were highest under A-5, resulting in increased exploitation and, consequently, the lowest average cumulative payoff among all personality conditions.
Figure 3: Average cooperation rates (left) and average cumulative payoffs (right) for GPT-4o under extreme personality-score manipulations. The top, middle, and bottom rows are defined in the same way as in Fig. 2.
For GPT-5, agreeableness similarly had the largest effect on cooperation, while manipulation of other personality traits produces little changed in behavior. Under the A-1 condition, cooperation rates fell to 0% across all opponent strategies, mirroring the behavior observed for GPT-4o. Under the A-5 condition, cooperation rates increased across strategies; however, the increase in cooperation against ALLD and RANDOM was smaller than that observed for GPT-4o. As a result, GPT-5 experienced less exploitation under high agreeableness than GPT-4o.
Figure 4: Average cooperation rates (left) and average cumulative payoffs (right) for GPT-5 under extreme personality-score manipulations. The top, middle, and bottom rows are defined in the same way as in Fig. 2.
Across all models, manipulation of agreeableness had the strongest and most consistent effect on cooperative behavior, whereas manipulation of other personality traits had comparatively limited influence. Moreover, as model generations progress from GPT-3.5-turbo to GPT-4o and GPT-5, the increase in cooperation against exploitative strategies under high agreeableness became smaller. This pattern indicates that newer models are better able to maintain restraint against non-cooperative opponents even when personality steering promotes cooperativeness.
4 Conclusion
In this study, we investigated how artificially controlled personality traits influence cooperative behavior in large language model agents. Using the Big Five framework, we first quantitatively measured basic personality profiles of multiple LLMs, and then examined their behavior in repeated Prisoner’s Dilemma games under conditions with and without explicit personality information. We further analyzed the effects of independently manipulating each personality dimension to extreme values.
Our results showed that agreeableness consistently had the strongest influence on cooperative behavior across all models, whereas manipulation of other personality traits produced comparatively limited effects. At the same time, explicitly providing personality information did not lead to identical behavior across models or conditions. Even when personality traits were controlled, LLM agents adjusted their decisions based on opponent strategies and interaction context.
Moreover, clear differences were observed across model generations. Earlier-generation models tended to exhibit increased cooperation accompanied by higher vulnerability to exploitation, while later-generation models showed more selective cooperation, particularly against exploitative opponents. These findings suggest that the impact of personality steering depends not only on the assigned personality traits but also on the strategic reasoning capabilities of the model.
Taken together, our study indicates that personality steering alone is insufficient to fully control cooperative behavior in LLM agents. Rather, cooperative outcomes emerge from the interaction between personality cues and model-specific decision-making processes. This highlights the importance of considering both personality and strategic reasoning when analyzing and designing cooperative AI agents.